Treatment 1

Treatment 2

Treatment 3

0

1

6

1

4

5

0

1

6

3

2

3

T = 4

T = 8

T = 20

SS = 6

SS = 6

SS = 6

   N = 12     G = 32 ƩX2= 138

1a.

Conduct a single-factor independent-measures ANOVA to test the hypothesis that there are significant differences in the mean scores among the three treatment conditions. Use α = .01.

The alternative hypothesis is

Group of answer choices

a) H1: μ1 ≠ μ2≠ μ3

b) H1: μ1 = μ2= μ3

c) all pairs of means are significantly different from each other

d) There is at least one significant mean difference: At least two means significantly differ from each other

1b.

The Critical F-value is:

Note: Leave the two decimal places after the decimal point

1c).

The F-statistic is

Note: If it is a decimal number with two or more than two places, leave only two decimal places after the decimal point and do not round.

1d)

Your decision is

Group of answer choices

a) Fail to reject the null hypothesis and conclude that there there is at least one significant mean difference

b) Fail to reject the null hypothesis and conclude that there are no significant differences among the three treatment conditions

c) Reject the null hypothesis and conclude that there is at least one significant mean difference

d) Reject the null hypothesis and conclude that there are no significant differences in mean scores among the treatment conditions

1e)

Calculate η2 (eta squared) as a measure of the effect size

η2 is

Note: If it is a decimal number with two or more than two places, leave only two decimal places after the decimal point and do not round.
Answer #1

Applying one way ANOVA from excel: data -data analysis:

Source of Variation SS df MS F P-value F crit
Between Groups 34.6667 2 17.3333 8.6667 0.0080 8.0215
Within Groups 18.0000 9 2.0000
Total 52.6667 11

The alternative hypothesis is d) There is at least one significant mean difference:

1b)

The Critical F-value is =8.02

1c). The F-statistic is =8.67

1d) c) Reject the null hypothesis and conclude that there is at least one significant mean difference
1e)

η2 =SSR/SST =34.6667/52.6667 =0.6582